# Engagement Maximization

### v4

@@ Abstract
We study a Bayesian agent receiving signals over time and then acting. The agent chooses when to stop and act, and prefers to act earlier all else equal. The signals are determined by a principal, who maximizes engagement (the total attention paid by the agent). We show that engagement maximization minimizes the agent’s welfare, induces excessive information acquisition (relative to an agent-optimal benchmark), and leads to extreme beliefs. The principal optimally sends only “suspensive signals” that lead the agent to become “less certain than the prior” and “decisive signals” that lead the agent to stop immediately.
@@ 1 Introduction
Our main result is the characterization of an optimal strategy for the principal. First, we completely characterize the optimal “overall” information structure (i.e. the signal-state joint distribution that is unconditional on time). It is characterized by the solution to an augmented static rational inattention (RI) problem: the information structure maximizes a linear combination of the instrumental value of information and the measure of information with endogenous weights. Second, we identify one optimal sequential information structure for the principal: the principal sends a “dilution” of the overall information structure (a compound Poisson process such that a signal arrives at a Poisson rate and upon arrival, the signal is distributed according to the overall information structure).
@@
Moreover, we show that any optimal strategy has the following features:
  • Agent welfare minimization: Under the principal’s optimally chosen information structure and the agent’s optimal stopping policy, the agent engages substantially with the platform but is no better off than if she did not use the platform and made her decision without acquiring any information. That is, engagement maximization by the principal minimizes the welfare of the agent. The agent’s ability to stop using the platform at any time does not allow her to extract any surplus. This result depends critically on the principal’s ability to send arbitrary signals to the agent; we provide an example in which the principal can choose from a more limited set of signals and show that such limits can allow the agent to extract positive surplus.
  • Belief polarization: When the agent chooses to stop and act after engaging with the platform, the agent will hold extreme beliefs, relative to a benchmark model in which the agent can choose both what information to receive and when to stop and act.
  • Learning paths: An optimal strategy contains (up to) two types of signals almost surely. The first type are decisive signals that are so informative that the agent makes a decision immediately after receiving the signal. The second type are suspensive signals that causes the agent’s beliefs to move to a posterior that is “more uncertain than the prior”.
@@
After presenting our main results of welfare-minimization and belief polarization, we consider several extensions of our baseline model. First, we consider the possibility that the agent does not fully attend to the signals sent by the principal. We show that it is without loss of generality to assume the agent pays full attention to the signals provided. Second, we show that the usual Bayesian persuasion assumption of commitment to a signal structure is without loss in our setting. Third, we consider the case in which there is an increasing opportunity cost of time and/or decreasing marginal utility from using the platform. This case is less tractable; to make progress, we impose additional symmetry assumptions, and derive conditions under which our main results continue to hold. Fourth, we discuss the importance of the role of jumps in beliefs in our model. The optimal belief process in our baseline model will involve rare but informative signal realizations that cause beliefs to jump. This result speaks to the sensational nature of the information that the principal optimally provides the agent. To clarify this point, we extend our analysis to the case in which the size of jumps in beliefs is arbitrarily restricted. We characterize how the restriction on jump size restricts the set of implementable strategies, and show that under certain circumstances, if beliefs must follow a continuous process, the principal and agent are in fact fully aligned and it is instead the agent who receives the full surplus. Fifth, we consider a modification of the model in which the user has unlimited information processing capacity and the platform maximizes time spent (as opposed to attention). We show that some of our results continue to apply to this modified model. Lastly, we apply our framework to a seemingly quite different setting, in which a teacher seeks to maximize the engagement of a student who cares only about passing a test. [^2] We are grateful to Emir Kamenica for suggesting this alternative setting as an application of our model.
1.1 Example
We illustrate the main model and the logic of our main result in a simple example. An agent faces a choice between two actions, $A=\{l,r\}$. The payoffs from the actions are uncertain and depend on the state of the world $x\in X=\{L,R\}$. The agent assigns equal prior probability to both states. The agent gets utility one when the chosen action matches the state ($l|L$ or $r|R$) and utility negative one otherwise ($l|R$ or $r|L$). The agent is impatient and pays a constant cost of delay of two utils per unit of time.
Information & capacity constraint:
For this example, we restrict attention to the following type of signals. The signal has two possible realizations $\{s_{l},s_{r}\}$, which arrives at Poisson rates that depend on the state (symmetrically):
L R
$s_{l}$ $\alpha\cdot p$ $\alpha\cdot(1-p)$
$s_{r}$ $\alpha\cdot(1-p)$ $\alpha\cdot p$
We assume that the agent is limited in her information processing capacity such that $\alpha$ and $p$ must satisfy
where $H^{S}(\cdot)$ is Shannon’s entropy. Note that $H^{S}\left(\frac{1}{2}\right)-H^{S}(p)$ is an increasing and convex function of $p$. Intuitively, both more frequent and more precise signals require processing “more information.” This constraint, which is a specialized version of the standard mutual information constraint used in the rational inattention literature, creates a trade-off between the frequency and the precision of signals. Formally, the constraint is equivalent to requiring that the mutual information of the state and signals received up to any time $t+\delta$, conditional on the signal history up to time $t$, is bounded by $\delta$.
The agent-optimal strategy:
We now explain why the agent is willing to wait for the arrival of the signal. If the agent has not observed either signal by time $t$, her beliefs will remain equal to her prior (and hence assign equal probability to each state). As a result, if the agent were to stop and act before observing either of the Poisson signals, her expected payoff would be zero. Because the problem is stationary, $V^{B}$ is also the agent’s expected utility at any time $t>0$, conditional on not having received either Poisson signal yet; it follows that the agent will strictly prefer waiting for the arrival of one of these signals to stopping. In fact, the results of Hébert and Woodford (2023) show that this signal structure and stopping strategy is optimal for the agent, both within the class of signals considered in this example and within the larger class of signals satisfying the mutual information bound.
The principal’s payoff depends on the total amount of information acquired by the agent, which we refer to as engagement. For the purpose of this example, we use Shannon’s entropy to both measure engagement and govern the rate at which the agent can process information. Our more general analysis will consider cases in which the principal’s engagement measure and the agent’s processing constraints differ. Under the proposed strategy, because the information capacity constraint always binds, the expected engagement is $\alpha^{-1}$, the expected decision time.
A better (but not optimal) strategy for the principal:
Let us now consider whether there are signals that are better for the principal, exactly satisfy the agent’s information processing constraint, and which the agent is willing to receive. Because such signals exactly satisfy the information processing constraint, maximizing engagement requires increasing the expected decision time. Within the class of Poisson signals, this is only possible if the signals become less frequent but more informative. Such signals are less desirable for the agent than the agent-optimal signals (by definition); but as established above, the agent strictly benefits from receiving the agent-optimal signals, and therefore might be willing receive somewhat less desirable signals as well.
Using these less frequent, more informative signals allows the principal to increase expected engagement ($\tilde{\alpha}^{-1}>\alpha^{-1}$) at the agent’s expense. It is straightforward in this example to see that using an even less frequent, even more informative signal would allow the principal to drive the agent to her indifference point (welfare minimization) while further increasing expected engagement. Our results will demonstrate that this is in fact an optimal strategy within a much more general class of sequential information structures. Note that utility is not transferable in this example (or in our general model); thus, the necessity of welfare minimization under the principal’s optimal strategy is not immediate from the setup of the model, and in fact will not hold when the principal is restricted to a subset of strategies (see Section 5).
@@ 1.2 Related Literature
Closely related to our work is the literature on dynamic Bayesian persuasion (e.g. Ely (2017); Renault et al. (2017); Ely and Szydlowski (2020); Orlov et al. (2020); Che et al. (2020)) that build on the static model of Kamenica and Gentzkow (2011). Our setting differs from most of these papers in that our principal’s payoff solely depends on the total attention paid by the agent, while the agent obtains pure instrumental value of information from solving a general decision problem. [^3] Our model is equivalent to one where the principal’s only goal is to ensure the agent continues to pay attention, as in Kawamura and Le Quement (2019). However, our principal presents unbiased information, and hence is not engaged in cheap talk (Crawford and Sobel (1982); Cheng and Hsiaw (2022)). Most of these papers focus on the “Bayesian persuasion” settings where the principal’s payoffs directly depend on the agent’s action and the maximization of “engagement” is a side effect. Two exceptions are Knoepfle (2020)and Koh and Sanguanmoo (2022), who also focus on maximizing “engagement” or “attention,” but without an information processing constraint. [^4] Knoepfle (2020)’s main focus is on the competition between multiple senders. The single sender case of Knoepfle (2020) is the closest to ours In both papers, the lack of an information constraint leads to the optimal persuasion strategy fully revealing the state. Our analysis emphasizes the beliefs the agent will arrive at, and how these differ between the principal’s optimal strategy and an agent-optimal benchmark, a comparison that is not possible absent information processing constraints. We elaborate further on the connection between our model and these models in Section 5.5. We assume both players are long-lived and have commitment power, in contrast to the limited commitment settings in Orlov et al. (2020); Che et al. (2020) or the myopic agent settings in Ely (2017); Renault et al. (2017). Interestingly, unlike the findings in these papers, neither commitment power nor forward-lookingness of the agent is necessary for sustaining our equilibrium strategy (see detailed discussions in Section 4).
@@
The rest of the paper is organized as follows. We begin in section 2 by describing the basic environment of our model. Section 3 characterizes optimal policy in our baseline model. Section 4 discusses the key implications, welfare minimization and belief polarization. Section 5 discusses several extensions of our baseline model, and in Section 6 we conclude.
@@ 2.1 The Agent’s Problem
At each time $t$, the agent can either stop and choose an action from a finite set $A$, or continue to acquire information. Let $\tau$ denote the time at which the agent stops and makes a decision, with $\tau=0$ corresponding to making a decision without acquiring any information. The agent receives utility $u_{a,x}$ if she takes action $a$ and the true state of the world is $x$, and pays a flow cost of delay per unit time, $\bar{\kappa}>0$, until an action is taken. This flow cost captures the opportunity cost of the agent’s time net of whatever utility is gained using the principal’s platform. In 5.3, we provide an extension where there is an increasing time-dependent flow cost of delay $\kappa(t).$ [^6] Note that we assume that the agent does not discount the utility from acting. Discounting, in our context, has the effect of changing the relative values of acting and never acting, and for this reason would complicate our analysis; see the more extensive discussions in Hébert and Woodford (2023) and Zhong (2022) on this point.
@@
The agent’s beliefs, $q_{t}$, will evolve as a martingale. This property follows from “Bayes-consistency.” In a single-period model, Bayes-consistency requires that the expectation of the posterior beliefs be equal to the prior beliefs. The continuous-time analog of this requirement is that beliefs must be a martingale.
@@ 2.2 The Principal’s Problem
The principal chooses the information the agent receives so as to maximize engagement. In the background, we have in mind that the principal is showing ads (unrelated to the agent’s current decision problem) or otherwise profiting [^7] Network effects, in particular the possibility that this agent’s attention increases utility other agents receive from the platform and thus the profits the principal can extract from those agents, are one relevant engagement maximization motive. from each additional increment of attention the agent pays to the information the principal provides.
We begin with defining a measure for the agent’s information processing using a continuous time version of what Caplin et al. (2022) call a “uniformly posterior-separable” cost function, as described in Hébert and Woodford (2023). Uniformly posterior-separable cost functions are defined in terms of a “generalized entropy function,” $H:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$. We assume that $H$ is strongly convex and twice continuously-differentiable. Then, define
@@
Here, $I_{t}^{H}$ denotes the cumulative information acquired by the agent measured by $H$. We assume the agent can process information at a rate defined by the constraint
@@
The principal’s goal, on the other hand, is to maximize “engagement.” Specifically, we assume that the principal earns profits in proportion to the information acquired by the agent, as measured by the generalized entropy function $G:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$ (a convex and twice continuously-differentiable function). Note that the function $G$ is not necessarily equal to the function $H$. The former determines the information that the principal would like the agent to acquire; the latter governs the rate at which the agent can acquire different kinds of information.
The principal’s goal is to design the agent’s belief process so as to maximize profits, taking into account the fact the agent will know the nature of the beliefs process and optimally choose when to stop paying attention. Let $\bar{q}_{0}\in\mathcal{P}(X)$ be the agent’s prior. The principal chooses his policies from the set $\mathcal{A}(\bar{q}_{0})$, which is the set of probability measures on $(\Omega,\mathcal{F})$ such that $q$ is martingale belief processes with $q_{0}=\bar{q}_{0}$, $I^{H}$ is an increasing adapted process with $I_{0}^{H}=0$, and non-negative stopping times $\tau$, such that (1) and (2) are satisfied for all $t\in[0,\tau)$. Formally, the principal chooses the probability measure $P$, which is equivalent to choosing the law of of the belief process $q$. In this sense, the principal can choose any càdlàg martingale belief process with $q_{0}=\bar{q}_{0}$, subject to the constraint imposed by the agent’s information processing capacity. [^9] An alternative interpretation is that the principal can send more information than the agent can process, in which case the agent chooses which information to attend to. Allowing the agent this choice cannot benefit the principal, and it is therefore without loss of generality to suppose the principal chooses a process that satisfies the agent’s information processing constraint.
@@
$J(\bar{q}_{0})$ $=\sup_{(P,I^{H},\tau)\in\mathcal{A}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$
@@
$\tau\in\arg\max_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\hat{u}(q_{\tau^{\prime}})-\bar{\kappa}\tau^{\prime}|\mathcal{F}_{0}].$

### v5

@@ Abstract
We investigate the management of information provision to maximize user engagement. A principal sequentially reveals signals to an agent who has a limited amount of information processing capacity and can choose to exit at any time. We identify a “dilution” strategy—sending rare but highly informative signals—that maximizes user engagement. The platform’s engagement metric shapes the direction and magnitude of biases in provided information relative to a user-optimal benchmark. Even without intertemporal commitment, the platform replicates full-commitment revenue by inducing the user’s belief to remain “as uncertain as” the prior until the rare, decisive signal arrives and induces stopping. We apply our results to two contexts: an ad-supported internet media platform and a teacher attempting to engage test-motivated students.
@@ 1 Introduction
A second and more subtle challenge is to manage the algorithm’s own “incentives”. Platforms often design these algorithms with the goal of providing reliable and wide-ranging content. However, subsequent learning based on user interactions can lead algorithms to maximize engagement by leading users down niche “rabbit holes ” (WSJ (2021)). This can lead to short term engagement gains but in the long-run will cause users to question the credibility of the platform. It is not obvious ex-ante whether concerns about credibility meaningfully constrain the algorithm design.
@@
Two modeling features allow us to derive the rich implications in this environment. First, our model allows the measures of (i) engagement, (ii) information processing capacity, and (iii) the value of information to differ. This allows our model to capture a rich set of possible incentive misalignments between the principle and the agent. Second, we analyze our model both with and without intertemporal commitment, which allows us to characterize the implications of a requirement for dynamic consistency on the dynamics of information provision.
@@
Our main result is the characterization of an optimal strategy for the principal. First, we completely characterize the optimal “overall” information structure (i.e. the signal-state joint distribution that describes the beliefs the agent will hold when choosing to stop). It is characterized by the solution to an augmented static rational inattention (RI) problem: the information structure maximizes a linear combination of the instrumental value, the engagement measure, and the informativeness measure of information with endogenous weights. Second, we identify one optimal sequential information structure for the principal: the principal sends a “dilution” of the overall information structure (a compound Poisson process such that a signal arrives at a Poisson rate and upon arrival, the signal is distributed according to the overall information structure).
@@
Having described a single optimal strategy, we then show that the following features must hold for any optimal strategy:
  • Engagement-driven biases: We compare the optimal information structure with the benchmark solution when the agent chooses both the information and the stopping time. We show that engaging with the platform leads to the agent being biased towards “more extreme” beliefs relative to the agent-optimal benchmark. More specifically, we show that the direction and magnitude of such biases are determined by the principal’s engagement measure.
    Leading Case: Revenue from all engagement. Suppose that the principal’s engagement measure is identical to the informativeness measure. We interpret this case as one in which the agent is exposed to ads in proportion to the context she receives, with no measurement of whether she attends to the ads. In this case, the optimal information structure is narrow but deep: the principal only provides information about the agent’s interested dimensions. This information is provided in a rare but chunky way, leading the agent to overspend their time and overdose information relative to the agent-optimal benchmark. This result rationalizes the commonly criticized rabbit-holing and polarization behavior of the recommendation algorithms when they target the maximization of engagement metrics.
    Leading Case: Revenue from decision-irrelevant engagement. Now suppose that the principal’s revenue comes from engagement with information that is irrelevant for decision. We interpret this case as one in which the agent is exposed to decision-irrelevant ads, and the principal benefits only when the agent engages with these ads (e.g. if ad revenue comes from the click of an ad.). In this case, we show that the optimal information structure is broad but noisy: the principal provides excess information (e.g. sponsored content) that the agent views as pure noise, intermixed with just enough decision-relevant information to keep the agent engaged. This result rationalizes the click-baiting and stealth marketing behavior of the recommendation algorithms when they target the maximization of clicks.
  • Credibility-driven dynamics: To study policies that are credible under limited intertemporal commitment, we introduce a notion of subgame perfection to our model. Firstly, we show that commitment has no value at all: the principal can achieve the same payoff from the optimal commitment policy even under limited commitment. However, limited commitment has strong implication on the dynamics of the policy: an optimal policy corresponds to a subgame perfect equilibrium if (and nearly only if) the induced belief process of the agent jumps between a small set of very special beliefs — beliefs at which the agent is “as uncertain as” the prior belief. As an implication, the optimal and credible signal process must be a generalized “dilution”: belief stays constant until a signal arrives at a Poisson rate. The signal brings the posterior belief either to the stopping region or to another such special interim belief. Consequently, the “dilution” policy we constructed earlier is indeed credible, and is the uniquely credible policy in certain symmetric settings.
    The intuition for the result is simple: We show that the principal-optimal policy leaves the agent with no surplus at the prior. If the agent were to become “more certain than” the prior at an interim stage, the surplus from future information would be negative, and the agent would stop engaging with the principal. On the other hand, if there agent were to become “less certain than” the prior at an interim stage, her surplus from future information would be positive. At this point, the principal would be tempted to have the algorithm adjust and extract more surplus, at the agents’ expense. Thus, the only possible interim beliefs are those that are exactly “as certain as” the prior belief. Our main result in Section 4 is a formalization of this idea, along with a precise definition of what it means to be more certain, less certain, or as certain as the prior.
@@
After presenting our main results , we consider several extensions of our baseline model. First, we consider the possibility that the agent does not fully attend to the signals sent by the principal. We show that it is without loss of generality to assume the agent pays full attention to the signals provided. Second, we show that the usual Bayesian persuasion assumption of commitment to a signal structure is without loss in our setting. Third, we consider the case in which there is an increasing opportunity cost of time and/or decreasing marginal utility from using the platform. This case is less tractable; to make progress, we impose additional symmetry assumptions, and derive conditions under which our main results continue to hold. Fourth, we discuss the importance of the role of jumps in beliefs in our model. The optimal belief process in our baseline model will involve rare but informative signal realizations that cause beliefs to jump. This result speaks to the sensational nature of the information that the principal optimally provides the agent. To clarify this point, we extend our analysis to the case in which the size of jumps in beliefs is arbitrarily restricted. We characterize how the restriction on jump size restricts the set of implementable strategies, and show that under certain circumstances, if beliefs must follow a continuous process, the principal and agent are in fact fully aligned and it is instead the agent who receives the full surplus. Fifth, we consider a modification of the model in which the user has unlimited information processing capacity and the platform maximizes time spent (as opposed to attention). We show that some of our results continue to apply to this modified model.
Throughout the paper, we discuss the example an ad-supported internet platform engaging a user. However, we pair this example with a second example, that of a teacher who seeks to maximize the engagement of a student who cares only about passing a test. [^2] We are grateful to Emir Kamenica for suggesting this alternative setting as an application of our model. Our theoretical framework applies to this setting without modification, and we discuss both examples to emphasize that engagement maximization is relevant in a wide range of contexts. As an example, the distinction we make between per-impression and per-click advertising in the context of an internet platform maps, in the teacher-student context, to the degree to which the teacher values test-relevant vs. test-irrelevant information. A teacher who values student learning but not test performance will provide just enough test-specific information to engage a student who cares only about the test, while providing additional content that is not part of the test. In contrast, a teacher who values all learning, regardless of whether it is covered on a test, will optimally provide only test-relevant information, while forcing students to learn more test-relevant information than they would choose to learn on their own.
@@ 1.1 Related Literature
Closely related to our work is the literature on dynamic Bayesian persuasion (e.g. Ely (2017); Renault et al. (2017); Ely and Szydlowski (2020); Orlov et al. (2020); Che et al. (2020)) that build on the static model of Kamenica and Gentzkow (2011). Our setting differs from most of these papers in that our principal’s payoff solely depends on the total attention paid by the agent, while the agent obtains pure instrumental value of information from solving a general decision problem. [^3] Our model is equivalent to one where the principal’s only goal is to ensure the agent continues to pay attention, as in Kawamura and Le Quement (2019). However, our principal presents unbiased information, and hence is not engaged in cheap talk (Crawford and Sobel (1982); Cheng and Hsiaw (2022)). Most of these papers focus on the “Bayesian persuasion” settings where the principal’s payoffs directly depend on the agent’s action and the maximization of “engagement” is a side effect. Two exceptions are Knoepfle (2020)and Koh and Sanguanmoo (2022), who also focus on maximizing “engagement” or “attention,” but without an information processing constraint. [^4] Knoepfle (2020)’s main focus is on the competition between multiple senders. The single sender case of Knoepfle (2020) is the closest to ours. In both papers, the lack of an information constraint leads to the optimal persuasion strategy fully revealing the state. Our analysis emphasizes the beliefs the agent will arrive at, and how these differ between the principal’s optimal strategy and an agent-optimal benchmark, a comparison that is not possible absent information processing constraints. We elaborate further on the connection between our model and these models in Section 5.5. We assume both players are long-lived and have commitment power, in contrast to the limited commitment settings in Orlov et al. (2020); Che et al. (2020) or the myopic agent settings in Ely (2017); Renault et al. (2017). Interestingly, unlike the findings in these papers, neither commitment power nor forward-lookingness of the agent is necessary for sustaining our equilibrium strategy (see Section 4.2).
@@
The rest of the paper is organized as follows. We begin in section 2 by describing the basic environment of our model. Section 3 characterizes optimal policy in our baseline model, with an emphasis on the beliefs the agent will hold when stopping. Section 4 discusses the dynamics of beliefs and the role of commitment. Section 5 discusses several extensions of our baseline model, and in Section 6 we conclude.
@@ 2.1 The Agent’s Problem
At each time $t$, the agent can either stop and choose an action from a finite set $A$, or continue to acquire information. Let $\tau$ denote the time at which the agent stops and makes a decision, with $\tau=0$ corresponding to making a decision without acquiring any information. The agent receives utility $u_{a,x}$ if she takes action $a$ and the true state of the world is $x$, and pays a flow cost of delay per unit time, $\bar{\kappa}>0$, until an action is taken. This flow cost captures the opportunity cost of the agent’s time net of whatever utility is gained as part of the information acquisition process (e.g. from using the principal’s platform). In 5.3, we provide an extension where there is an increasing time-dependent flow cost of delay $\kappa(t).$ [^6] Note that we assume that the agent does not discount the utility from acting. Discounting, in our context, has the effect of changing the relative values of acting and never acting, and for this reason would complicate our analysis; see the more extensive discussions in Hébert and Woodford (2023) and Zhong (2022) on this point.
@@
The agent’s beliefs, $q_{t}$, will evolve as a martingale. This property follows from “Bayes-consistency.” In a single-period model, Bayes-consistency requires that the expectation of the posterior beliefs be equal to the prior beliefs. The continuous-time analog of this requirement is that the belief process is a martingale.
@@ 2.2 The Principal’s Problem
The principal chooses the information the agent receives so as to maximize engagement. In the background, we have in mind that the principal is showing ads or otherwise profiting [^7] Network effects, in particular the possibility that this agent’s attention increases utility other agents receive from the platform and thus the profits the principal can extract from those agents, are one relevant engagement maximization motive. from each additional increment of attention the agent pays to the information the principal provides.
We begin with defining a measure for the agent’s information processing using a continuous time version of what Caplin et al. (2022) call a “uniformly posterior-separable” cost function, as described in Hébert and Woodford (2023). Uniformly posterior-separable cost functions are defined in terms of a “generalized entropy function,” $H:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$. We assume that $H$ is strongly convex and twice continuously-differentiable. Using $H$, we define the cumulative information acquired, $I^{H}_{t}$, by
@@
We assume the agent can process information at a rate defined by the constraint
@@
The principal’s goal is to maximize “engagement,” which is not necessarily the same as maximizing the information acquired. Specifically, we assume that the principal earns profits in proportion to cumulative information acquisition as measured by the generalized entropy function $G:\mathcal{P}(X)\rightarrow\mathbb{R}_{+}$ (another convex and twice continuously-differentiable function). The function $G$ is not necessarily equal to the function $H$, although this will be one case of interest. The former determines the information that the principal would like the agent to acquire; the latter governs the rate at which the agent can acquire different kinds of information. We will assume that there is always some information the principal would like the agent to acquire (i.e. that the Hessian of $G$ is non-zero on the interior of the simplex).
The principal’s goal is to design the agent’s belief process so as to maximize profits, taking into account the fact the agent will know the nature of the beliefs process and optimally choose when to stop paying attention. Let $\bar{q}_{0}\in\mathcal{P}(X)$ be the agent’s prior. The principal chooses his policies from the set $\mathcal{A}(\bar{q}_{0})$, which is the set of probability measures on $(\Omega,\mathcal{F})$ such that $q$ is martingale belief processes with $q_{0}=\bar{q}_{0}$, increasing adapted processes $I^{H}$ with $I_{0}^{H}=0$, and non-negative stopping times $\tau$, such that (1) and (2) are satisfied for all $t\in[0,\tau)$. Formally, the principal chooses the probability measure $P$, which is equivalent to choosing the law of of the belief process $q$. In this sense, the principal can choose any càdlàg martingale belief process with $q_{0}=\bar{q}_{0}$, subject to the constraint imposed by the agent’s information processing capacity. [^9] An alternative interpretation is that the principal can send more information than the agent can process, in which case the agent chooses which information to attend to. Allowing the agent this choice cannot benefit the principal, and it is therefore without loss of generality to suppose the principal chooses a process that satisfies the agent’s information processing constraint.
@@
$J(\bar{q}_{0})$ $=\sup_{(P,I^{H},\tau)\in\mathcal{A}(\bar{q}_{0})}\mathbb{E}^{P}[G(q_{\tau})-G(\bar{q}_{0})|\mathcal{F}_{0}]$ (P)
@@
$\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|\mathcal{F}_{0}]\geq\sup_{\tau^{\prime}\in\mathcal{T}}\mathbb{E}^{P}[\hat{u}(q_{\tau^{\prime}})-\bar{\kappa}\tau^{\prime}|\mathcal{F}_{0}].$